# Improving Genomic Prediction using High-dimensional Secondary Phenotypes: the Genetic Latent Factor Approach

### v1

@@ 6. Data availability
We used the keras (v2.13.0) and tensorflow (v2.15.0.9) R packages as an interface to Python (v3.10) and TensorFlow (v2.13.1) to implement the deep learning [9, 10].
S.6. Interpretability
@@
We have stressed the great ease of interpreting gfBLUP’s model parameters in terms of biological relevance. Here we present some additional results with regards to interpretability of the approaches that we compared gfBLUP to. We ran si-, lsBLUP, and MegaLMM on the non-redundancy filtered hyperspectral data from 10-03-2015 as we did for gfBLUP in Section 4.2. We then looked at the relevant model parameters that included the selection index weights for siBLUP, LASSO coefficients for lsBLUP, and (mean) posterior loadings for MegaLMM.
Figure S1. LASSO regression coefficients (A) and selection index weights (B) for the $62$ wavelengths obtained from running lsBLUP and siBLUP on the hyperspectral data from 10-03-2015. Estimated genetic correlations between the LASSO prediction (LS) or selection index (SI) and yield (Y) are also shown.
Figure S2. MegaLMM mean posterior loadings for the $62$ wavelengths on the $5$ (A) and $10$ (B) factors used to run MegaLMM on the hyperspectral data from 10-03-2015. Loadings of yield on the factors are also shown.
@@
Starting with lsBLUP and siBLUP (Figure S1) we immediately recognize the difficulties associated with compressing the collinear data to a single dimension. In the case of lsBLUP we see sign-flipping in the coefficients around the $700$nm wavelength. Reflectivities between $650$nm and $700$nm show an increasingly positive relationship with yield following from the genetic correlation of $0.81$ between the LASSO prediction and yield. After $700$nm the relationship rapidly becomes negative, before moving towards 00 again.
The selection index weights of siBLUP are difficult to interpret as well. The relationship between reflectivity and yield follows a somewhat similar pattern to lsBLUP, given the highly positive genetic correlation between the selection index and yield ($\rho^{g}=0.81$). The relationship becomes increasingly positive, followed by a strong decrease after $700$nm, before finally increasing again. The relationship then continues to oscillate in the remainder of the infra-red part of the spectrum.
@@
While si- and lsBLUP do not require any hyperparameter tuning, their lack of flexibility with regards to the selection of an appropriate number of dimensions clearly limits their interpretability in the case of more complex data.
@@
MegaLMM does provide the flexibility of choosing the number of factors ($M$) and thereby the dimensionality, but leaves this choice to the user. As a result, the ease of interpretation heavily relies on an appropriate choice of $M$, which may not be easy to make without any prior knowledge (Figure S2). We opted for values of $M=5$ and $M=10$, which seemed appropriate given the fact we found $3$ factors using gfBLUP. Note that while the main factors representing the infrared and visible part of the spectrum are still present (F5 and F2 for $M=5$, F3 and F1 for $M=10$), ease of interpretation goes down as the number of factors increases. This is mainly due to the loadings of yield, which are analogous to the genetic correlations between yield and factors in gfBLUP, becoming divided over several relevant factors as well as non-relevant factors.
@@
Furthermore, we observe strong non-stationary behavior in the loading MCMC chains for both $M=5$ (Figure S3) and $M=10$ (Figure S4), even after $10{,}000$ burn-in and a further $100{,}000$ posterior iterations for wavelengths around $700$nm (thinning rate of $2$).
@@
S.7. Computational complexity
Obtaining test set predictions for a single CV2 hyperspectral dataset (including all $620$ wavelength-date features) took approximately $1.53$ seconds for the univariate gBLUP model, $7.83$ seconds for lsBLUP, $218$ seconds for siBLUP, $43.3$ seconds for gfBLUP, $1030$ seconds for MegaLMM, and $93.2$ seconds for the CV2 multivariate MLP.
We also simulated datasets containing $p=100$ to $p=2500$ secondary features representing $10$ underlying factors. We analyzed these datasets using gfBLUP, obtaining five runtime measurements (after a single warmup run) for each step in the gfBLUP prediction pipeline at each value of $p$, showing that the regularization dominates the pipeline’s total runtime throughout the range of $p$ (Figure S5).
All computational times were obtained on Windows 11 Pro 23H2, R v4.2.2, using an Intel i5-13600KF CPU. Runtimes for deep learning were obtained using TensorFlow v2.15.1 and Python v3.10 through the keras v2.15.0 and tensorflow v2.16.0 R packages.
Figure S5. Composition of total runtime of gfBLUP at different numbers of secondary features $p$. The red line shows the total runtime in seconds.
S.8. CV2 using vegetative data
From a practical point of view, including test set secondary data measured during the full crop growth cycle (including grain filling) in CV2 single trial genomic prediction makes little sense. A more relevant scenario is to only include test set secondary features measured during the vegetative phase. Plots in field trials can then be culled based on multivariate genomic predictions to avoid crossing and harvesting of unpromising genotypes, saving a significant amount of resources.
Figure S6. Prediction accuracies for several methods in the CV1, CV2, and CV2VEG (secondary features measured during the vegetative phase only) scenarios for the wheat yield and hyperspectral reflectivities dataset. Redundancy filtering was carried out for all multivariate analyses with a threshold $\tau=0.95$.
Figure S6 provides results for such a scenario. While CV2 predictions using features from the vegetative phase only (CV2VEG) are lower than those obtained when including all data (CV2), multivariate approaches still outperform the univariate model by a significant margin.
S.9. Redundancy filtering
Figure S7 provides a graphical overview of the features that remained after filtering at a threshold of $\tau=0.95$.
Figure S7. Features (wavelength-date combinations) that remain after redundancy filtering using a threshold of $\tau=0.95$.

### v3

@@ 5. Discussion
We used the keras (v2.15.0), tensorflow (v2.16.0), and reticulate (v1.40.0) R packages as an interface to Python (v3.10.16) and TensorFlow (v2.15.1) to implement the deep learning (Chollet et al., 2015; Abadi et al., 2016).
S.5. Interpretability
@@
We have stressed the great ease of interpreting glfBLUP’s model parameters in terms of biological relevance. Here we present some additional results with regards to interpretability of the approaches that we compared glfBLUP to. We ran si-, lsBLUP, and MegaLMM on the non-redundancy filtered B5IR hyperspectral data from 10-03-2015 as we did for glfBLUP. We then looked at the relevant model parameters that included the selection index weights for siBLUP, LASSO coefficients for lsBLUP, and (mean) posterior loadings for MegaLMM.
@@
Starting with lsBLUP and siBLUP (Figure S1) we immediately recognize the difficulties associated with compressing the collinear data to a single dimension. In the case of lsBLUP we see sign-flipping in the coefficients around the $760$nm wavelength. Reflectivities around $660$nm, $720$nm, and $760$nm show a negative relationship with yield following from the genetic correlation of $0.81$ between the LASSO prediction and yield. After $760$nm the relationship rapidly becomes positive, before moving towards 00 again.
The selection index weights of siBLUP are difficult to interpret as well. The relationship between hyperspectral reflectivity and yield follows a highly irregular pattern along the spectrum.
@@
While si- and lsBLUP require minimal hyperparameter tuning, their lack of flexibility with regards to the selection of an appropriate number of dimensions clearly limits their interpretability in the case of more complex data.
@@
MegaLMM does provide the flexibility of choosing the number of factors ($M$) and thereby the dimensionality, but leaves this choice to the user. As a result, the ease of interpretation heavily relies on an appropriate choice of $M$, which may not be easy to make without any prior knowledge (Figure S2). We opted for values of $M=3$, $M=5$, and $M=10$, which seemed appropriate given the fact we found $3$ factors using glfBLUP. Note that while the main factors representing the infrared and visible part of the spectrum are still present (F1 and F2 for $M=3$, F1, F2, and F3 for $M=5$, F6, F3, F2, F5, and F4 for $M=10$), ease of interpretation quickly goes down as the number of factors increases. This is mainly due to the loadings of yield, which are analogous to the genetic correlations between yield and factors in glfBLUP, becoming divided over a large number of factors. We also stress that the choice to use a relatively low number of factors here was informed by the results obtained using glfBLUP. Runcie et al. (2021) use $100$ factors to analyze the hyperspectral data in their paper, which would likely make interpretation even more difficult.
@@
Furthermore, we observe non-stationary behavior in the loading MCMC chains for $M=3$ (Figure S3), $M=5$ (Figure S4), as well as $M=10$ (Figure S5 and S6), even after $10{,}000$ burn-in and a further $100{,}000$ posterior iterations for wavelengths around $700$nm (thinning rate of $2$). Non-stationarity is also visible in the traceplots for the loadings of yield on the different factors (Figure S7).
Figure S1. LASSO regression coefficients (A) and selection index weights (B) for the $62$ wavelengths obtained from running lsBLUP and siBLUP on the B5IR hyperspectral data from 10-03-2015. Estimated genetic correlations between the LASSO prediction (LS) or selection index (SI) and yield (Y) are also shown.
Figure S2. MegaLMM mean posterior loadings for the $62$ wavelengths on the $3$ (A), $5$ (B), and $10$ (C) factors used to run MegaLMM on the B5IR hyperspectral data from 10-03-2015. Mean posterior loadings of yield on the factors are also shown.
Figure S3. MegaLMM ($3$ factors, $100{,}000$ iterations with thinning rate of $2$ after $10{,}000$ burn-in iterations) traceplots for the posterior samples for loadings of wavelengths around 700 nm on the $3$ factors (top to bottom).
@@
Figure S7. MegaLMM posterior iterations for the loadings of yield on the factors for $M=3$ (A), $M=5$ (B), and $M=10$ (C) for the B5IR hyperspectral data from 10-03-2015.
S.6. Computational complexity
Obtaining test set predictions for a single CV2 hyperspectral dataset (including all $620$ wavelength-date features) took approximately $1.5$ seconds for the univariate gBLUP model, $7$ seconds for lsBLUP, $200$ seconds for siBLUP, $34$ seconds for glfBLUP, $1090$ seconds for MegaLMM, and $60$ seconds for the CV2 multivariate MLP.
We also simulated datasets containing $p=100$ to $p=2500$ secondary features representing $10$ underlying factors. We analyzed these datasets using glfBLUP, obtaining five runtime measurements (after a single warmup run) for each step in the glfBLUP prediction pipeline at each value of $p$, showing that the regularization dominates the pipeline’s total runtime throughout the range of $p$ (Figure S8).
All computational times were obtained on Windows 11 Pro 23H2, R v4.2.2, using an Intel i5-13600KF CPU.
Figure S8. Composition of total runtime of glfBLUP at different numbers of secondary features $p$. The red line shows the total runtime in seconds.
S.7. Redundancy filtering
Figures S9 and S10 provide graphical overviews of the features that remained after redundancy filtering at a threshold of $\tau=0.95$ for both managed treatments B5IR and HEAT.
Figure S9. Features (wavelength-date combinations) that remain after redundancy filtering using a threshold of $\tau=0.95$ for the B5IR treatment.
Figure S10. Features (wavelength-date combinations) that remain after redundancy filtering using a threshold of $\tau=0.95$ for the HEAT treatment.